OpenStudy (lasttccasey):

I need some help solving this real quick. Two equations two unknowns. Trig involved.

OpenStudy (lasttccasey):

Given:\[2F_B=F_A\]Solve:\[F_B*\cos(\theta)-F_A*\cos(30)=0\]\[F_B*\sin(\theta)+F_A*\sin(30)-98.1=0\]

OpenStudy (anonymous):

\[ F_B\cos(\theta)-2F_B\cos(30)=0 \]\[ \cos(\theta)= 2\cos(30) \]

OpenStudy (lasttccasey):

The answer to θ is 64.3° according to the book. Any ideas on getting that?

OpenStudy (anonymous):

Your equations must be wrong.

OpenStudy (anonymous):

It's not possible for \(F_B=2F_A\) if \[ F_B\cos(\theta)-2F_B\cos(30^\circ)=0 \]